# Comparing Bayesian and Frequentist Inference in Biological Models: A Comparative Analysis of Accuracy, Uncertainty, and Identifiability

### v1

@@ 6.3 GLM model for mpox epidemic in the USA
In this section, we present the result of fitting the generalized logistic model (GLM) to the 2022 U.S. mpox epidemic.
Figure 14 shows GLM fits to U.S. mpox incidence; both frequentist (QDF) and bayesian (BFF) capture the rapid rise and decline. The first column represents the results obtained using the QDF, while the second column shows the results from the BFF method.
Figure 14: GLM fits to U.S. mpox incidence using frequentist inference (QDF) and bayesian inference (BFF); both reproduce the observed rise–fall pattern.
Table 20 reports parameter estimates $(r,p,k)$ for the mpox GLM; the methods give similar central values with modest interval differences.
@@
Table 20: Parameter estimates for the GLM (mpox): both frameworks yield similar point estimates for $r$, $p$, and $k$ with modest differences in uncertainty.
Table 21 compares MAE, MSE, WIS, and 95% PI; QDF attains lower error metrics with comparable coverage.
@@
Table 21: Performance metrics for the GLM (mpox): QDF achieves lower MAE, MSE, and WIS with coverage near 95% PI, comparable to BFF.
Table 22 summarizes the convergence for the Bayesian fit, with all $\hat{R}=1$ indicating good mixing.
@@
$\hat{R}$ (BFF) 1 1 1
Table 22: Convergence diagnostics for the Bayesian GLM (mpox): all parameters have $\hat{R}=1$, indicating good MCMC convergence.
6.3.1 Comparison of the parameters estimation of GLM model using Mpox data
Figure 15 compares $(r,p,k)$ between methods, showing close agreement and slightly different uncertainty widths. Each point represents the posterior mean, with horizontal lines indicating the 95% PI. The comparison highlights the consistency in parameter estimation across methods while also revealing subtle differences in uncertainty ranges.
Figure 15: Parameter comparison for the GLM (mpox): $r$, $p$, and $k$ estimates are consistent across QDF and BFF, with slightly different interval widths.
6.3.2 Performance metrics comparison of MAE, MSE, WIS, and 95% PI for the GLM Model using mpox data
Figure 16 visualizes the error and coverage metrics. Both methods produced similar results for all four evaluation metrics: MAE, MSE, WIS, and 95% PI. The MAE values are nearly identical, indicating comparable point forecast accuracy. MSE is slightly lower for QDF, suggesting marginally better handling of larger errors, while The (WIS) values are very close, reflecting similar quality in predictive intervals. Both approaches achieved nearly identical 95% PI, indicating that their uncertainty quantification is equally well-calibrated. Overall, both methods performed comparably for the mpox data in the GLM model.
Figure 16: Performance comparison for the GLM (mpox): low errors and near-nominal 95% PI indicate good calibration on this dataset.
6.3.3 Posterior Histograms of GLM Parameters ($r,k,p$) for Mpox Data
The posterior (BFF) and sampling (QDF) histograms of the GLM parameters $(r,k,p)$ for the mpox incidence data are shown in Figures (S9)–(S10). Each panel summarizes the uncertainty with the mean, median, and 95% CI, allowing side-by-side comparison of BFF and QDF on the same dataset. Figures(S10) You will see $r$ concentrated roughly between $\sim 2$ and $\sim 3.3$, $p$ around $\sim 0.80$$0.85$, and $k$ centered near $\sim 3.0\times 10^{4}$ (tight unimodal shapes $\rightarrow$ with tight identifiability). Figures(S9) The printed stats on the panels show very tight posteriors (e.g., $p\approx 0.98,k\approx 2734,r\approx 0.33$). The GLM parameters are clearly identified under both methods; verify the Figure (S9) labeling in your source before submission.
6.3.4 GLM Parameter Settings for Mpox Data (BFF, Normal Error)
@@
6.3.5 GLM Parameter Settings for Mpox Data (QDF, Normal Error)
@@ 6.4 SEIUR model of COVID-19 cases in Spain during the first wave
In this section, we present the parameter estimation, performance metrics, and convergence analysis obtained by fitting the COVID-19 first-wave data from Spain to a compartmental SEIUR (Susceptible-Exposed-Infectious-Recovered model).
Figure 17 shows the SEIUR fits of the first wave in Spain; both frameworks track the rise and decline in incidence.
Figure 17: SEIUR fit to Spain’s first COVID-19 wave using QDF and BFF; both reproduce the observed rise–fall in incidence.
Table 25 reports $(\beta_{0},\beta_{1},q_{1},\rho,\kappa,\gamma_{1})$; estimates are comparable between methods with some differences in $\rho$ and $q_{1}$.
@@
Table 25: SEIUR parameter estimates $(\beta_{0},\beta_{1},q_{1},\rho,\kappa,\gamma_{1})$ for Spain first wave under BFF and QDF.
Table 26 compares MAE, MSE, WIS, and 95% PI; BFF attains lower errors and similar coverage than QDF.
@@
Table 26: SEIUR performance metrics (MAE, MSE, WIS, 95% PI): BFF achieves lower errors with similar coverage relative to QDF.
Table 27 lists convergence for the Bayesian fit; $\hat{R}$ values indicate acceptable mixing overall.
@@
$\hat{R}$ (BFF) 1.18 1.22 1.17 1 1.07 1.21
Table 27: Convergence diagnostics for the bayesian SEIUR fit; $\hat{R}$ values indicate acceptable chain mixing.
6.4.1 Comparison of the parameters estimation of SEIUR model
Figure 18 compares the parameters ($\beta_{0}$, $\beta_{1}$, $q_{1}$, $\rho$, $\kappa$ and $\gamma_{1}$ )estimates between methods, highlighting a larger uncertainty for several rates under BFF.
Figure 18: Parameter comparison for SEIUR (Spain): cross-method estimates for $(\beta_{0},\beta_{1},q_{1},\rho,\kappa,\gamma_{1})$ with differing uncertainty widths.
6.4.2 Performance metrics comparison of MAE, MSE, WIS, and 95% PI for the SEIUR model
Figure 19 visualizes MAE, MSE, WIS, and PI; BFF is best in all error metrics with similar coverage.
Figure 19: SEIUR performance comparison (Spain): BFF shows lower MAE/MSE/WIS with similar 95% PI.
6.4.3 Posterior Histograms of SEIUR Parameters ($\beta_{0},\beta_{1},q_{1},\rho,\kappa,\gamma_{1},N$)
Figures S11S12 show the distributions of the SEIR parameters under BFF and QDF. Each panel summarizes the uncertainty with the mean, median, and 95% CI) facilitating the method-by-method comparison. Figure S11 (BFF): Posteriors are moderately wide for several rates, indicating only partial identifiability from the data alone. Examples printed on the graphs: $\beta_{0}$ median $2.18$ with 95% CI $[1.60,3.09]$; $\beta_{1}$ median $1.69$ $[1.14,2.48]$; $q$ median $0.88$ $[0.64,1.41]$; $\rho$ median $0.60$ $[0.28,0.92]$; $\kappa$ median $2.39$ $[1.23,3.73]$; $\gamma$ median $1.78$ $[1.22,2.59]$. These wide intervals highlight the coupling among parameters and the limited information in the observed outputs. Figure S12 (QDF): Distributions appear more concentrated for several parameters (narrower support on the x-axis, unimodal), consistent with tighter point estimates under QDF. Use it in conjunction with Figure S11 to contrast where the BFF uncertainty bands are widest. BFF histograms are relatively broad for multiple rates, reflecting partial identifiability, while QDF histograms show sharper concentrations, indicating more stable estimates. This contrast underscores where additional data or reparameterization could improve identifiability.
6.4.4 Comparison of the parameters of the SEIR model (BFF) using a Normal error structure
@@
Table 28: Parameter ranges for the SEIUR model under BFF with Normal error.
6.4.5 Comparison of the parameters of the SEIUR model (QDF) using a Normal error structure
@@
Table 29: Parameter ranges for the SEIUR model under QDF with Normal error.
@@ 7 Discussion
This comparative analysis of Bayesian and Frequentist inference across ecological, clinical, and epidemiological systems reveals a clear practical distinction: Bayesian methods (BFF) deliver superior uncertainty quantification and diagnostic transparency, while Frequentist methods (QDF) often achieve lower point-error metrics with greater computational efficiency. The relative performance of each framework depends critically on the richness of the data, the complexity of the model, and the degree of structural identifiability. Our findings demonstrate that neither approach universally dominates; instead, the optimal choice depends on the specific characteristics of the modeling problem.
7.1 Model-Specific Performance
7.1.0.1 Lotka–Volterra Model.
@@ 7.1 Model-Specific Performance
For the LV model, the performance varied substantially in all three observation scenarios. When both prey and predator populations were observed (LV-1), both methods successfully recovered parameter values, with QDF achieving slightly lower point errors for predator dynamics (MAE 2.87 vs. 4.95) (Table 5) but BFF achieved perfect prediction interval coverage for prey (100% vs. 90.48%) (Table 5). Under partial observation with prey-only data (LV-3), QDF produced a sharper prediction of points , reducing MAE from 4.60 (BFF) to 3.30, MSE from 27.34 to 15.83 and WIS from 2.80 to 2.10 (Table 11). However, BFF maintained a superior calibration with 100% PI compared to QDF 90.48%. Convergence diagnostics confirmed stable MCMC sampling for BFF, with all LV parameters achieving $\hat{R}\approx 1.01$ (Table 6), indicating reliable posterior inference despite partial observability.
The structural identifiability analysis (Section 5) explains these patterns: full observation of both species enables unique parameter recovery, while partial observation creates identifiability deficits that neither inference framework can overcome without known initial conditions. The empirical results align with the SI predictions, demonstrating that observability fundamentally constrains inference quality regardless of the statistical method employed.
7.1.0.2 Generalized Logistic Model (Lung Injury and Mpox).
For GLM applied to lung injury and mpox outbreaks, QDF consistently outperformed BFF in point forecast accuracy while maintaining comparable or superior uncertainty calibration. In the mpox epidemic, QDF achieved lower MAE (9.45 vs. 11.74), MSE (150.47 vs. 399.43), and WIS (5.8 vs. 7.18), with perfect PI coverage (100% vs. 97.8%) (Table 21). BFF demonstrated excellent convergence ($\hat{R}=1$) (Table 17) and (Table 22) for all parameters $(r,p,K)$, confirming that the posterior was well-sampled despite the less favorable performance metrics. The lung injury results exhibited the same qualitative pattern, with QDF achieving lower point errors while maintaining well-calibrated prediction intervals.
The robust structural identifiability of the GLM (Section 5) explains why both methods succeeded in parameter recovery. The single-equation structure combined with direct observation of incident cases ensures that all parameters $(r,p,K)$ are uniquely identifiable under both known and unknown initial conditions. In such well-identified settings, the computational efficiency and optimization focus of QDF provide advantages for point estimation, while BFF’s probabilistic framework incurs additional computational cost without substantial gains in forecast accuracy.
7.1.0.3 SEIUR Model (COVID-19 in Spain).
In contrast to the GLM results, BFF outperformed QDF in all metrics for the SEIUR model: MAE 294.57 vs. 352.30, MSE 210,754 vs. 311,054, WIS 211.30 vs. 250.46, and PI coverage 94% vs. 92% (Table 26). This reversal of relative performance reflects the increased complexity of the model, latent compartments, and limited structural identifiability of the SEIUR framework. SI analysis (Section 5) demonstrates that when only incident reported cases are observed, most parameters and state variables ($S$, $E$, II, $U$, $R$, NN, $\rho$) are structurally unidentifiable when initial conditions are unknown. Only cumulative cases $C$ incubation rate $\kappa$, and recovery rate $\gamma_{1}$ remain identifiable.
In this challenging setting, Bayesian inference benefits from prior regularization that constrains the otherwise ill-posed inverse problem. The posterior distribution incorporates both the likelihood of the data and prior information, enabling more stable parameter estimation and forecast generation despite fundamental identifiability limitations. Frequentist optimization, on the contrary, struggles with the high-dimensional, poorly constrained parameter space, leading to less reliable point estimates and wider prediction intervals.
7.2 Role of Structural Identifiability
The structural identifiability analysis provides critical insight into why inference performance varies across models and observation scenarios. For the LV model, global identifiability requires observation of both prey and predator; partial observation creates parameter non-uniqueness that neither BFF nor QDF can resolve (Table 2). For the GLM, robust identifiability under all observation schemes ensures reliable inference for both frameworks. For the SEIUR model, severe identifiability deficits constrain parameter recovery regardless of the inference method, though Bayesian priors provide some regularization benefit.
These findings underscore that structural identifiability sets fundamental limits on what can be learned from the data. No statistical framework, no matter how sophisticated, can uniquely recover structurally unidentifiable parameters. SI analysis should therefore precede empirical modeling to clarify what questions the data can and cannot answer.
7.3 Practical Implications
Our results provide practical guidance for selecting between Bayesian and Frequentist inference frameworks:
7.3.0.1 Use Frequentist methods (QDF)
when: (1) data are abundant and of high quality, (2) models are structurally identifiable, (3) computational speed is critical, (4) the primary objective is minimizing point forecast error, and (5) model complexity is low to moderate. The GLM applications exemplify these conditions.
7.3.0.2 Use Bayesian methods (BFF)
when: (1) data are sparse, noisy, or partially observed, (2) models contain latent states or unobserved compartments, (3) structural identifiability is limited, (4) comprehensive uncertainty quantification is essential, (5) prior information is available to constrain inference, and (6) diagnostic assurance via convergence metrics is important. The SEIUR application exemplifies these conditions.
7.3.0.3 Improve both methods
by: (1) increasing observational coverage (e.g., measuring both prey and predator rather than one species), (2) obtaining accurate estimates of initial conditions, (3) incorporating auxiliary data sources to break parameter non-uniqueness, and (4) using SI-guided experimental design to ensure parameters of interest are identifiable.
When practical identifiability is limited or interval calibration is critical, BFF is advantageous because posterior sampling naturally captures parameter and predictive uncertainty while providing convergence assurance (e.g., GLM $\hat{R}=1$; LV $\hat{R}\approx 1.01$). When the priority is minimizing point forecast error and data are informative, QDF can be preferable (e.g., LV prey-only; GLM mpox), though prediction intervals may be under-calibrated relative to BFF (e.g., LV prey-only coverage 90.48% vs. 100%). Across all systems, adding informative observations improves both methods (LV predator-prey vs. single-series), and SI analysis clarifies when further accuracy gains are impossible without additional data or constraints.
7.4 Limitations and Future Directions
This study standardized model structures, observation schemes, and error distributions to isolate the effects of the inference framework. Although this design enables for clean comparison, several limitations warrant consideration. First, we assumed normal error structures for both methods; alternative distributions (e.g., negative binomial for count data) might shift relative performance. Second, we analyzed each model independently; hierarchical or multi-level structures could favor Bayesian inference more strongly. Third, our SI analysis assumed perfect, noise-free data; practical identifiability under realistic noise levels merits further investigation.
@@ 7.4 Limitations and Future Directions
Future research could explore hybrid inference strategies that combine the computational efficiency of optimization-based methods with the uncertainty quantification strengths of Bayesian sampling. The SI-guided experimental design could determine the optimal measurement strategies before data collection. Incorporating machine learning approaches such as neural ODEs giampiccolo2024hnode or physics-informed neural networks may offer complementary advantages. Extending the comparison to spatially explicit models, agent-based simulations, and stochastic epidemic frameworks would broaden the applicability of these findings.
8 Conclusion
Mathematical models based on ordinary differential equations are essential tools for understanding and predicting dynamics in ecological and epidemiological systems. The reliability of model-based inferences is critical to accurate parameter estimation and proper assessment of parameter identifiability. This study provides a systematic comparison of Bayesian and Frequentist inference frameworks across three biological models LV predator-prey dynamics, the generalized logistic model, and the SEIUR epidemic model applied to four real world datasets spanning ecology and epidemiology.
Our analysis demonstrates that neither Bayesian nor Frequentist methods universally outperform the other; instead, the optimal framework depends on data characteristics, model structure, and the degree of structural identifiability. Frequentist methods (implemented via QDF) excel when data are abundant, models are well identified, and computational efficiency is prioritized, as demonstrated by superior point forecast accuracy for the GLM applications to lung injury and mpox outbreaks. Bayesian methods (implemented via BFF) excel when data are sparse or noisy, models contain latent compartments, and comprehensive uncertainty quantification is essential, as demonstrated by superior performance for the SEIUR model of COVID-19 in Spain.
Structural identifiability analysis emerged as a critical prerequisite for meaningful inference. Our symbolic identifiability results using StructuralIdentifiability.jl revealed that observability fundamentally constrains parameter recovery: full observation of both predator and prey populations in the LV model enabled complete parameter identification, while partial observation created identifiability deficits that neither statistical framework could overcome without known initial conditions. Similarly, the SEIUR model’s limited identifiability when only reported cases are observed explains why both methods struggled with parameter uncertainty, although Bayesian priors provided some regularization benefit.
Integration of structural identifiability analysis with empirical inference provides a principled workflow for biological modeling. First, conduct a SI analysis to determine which parameters can theoretically be recovered from the planned observations. Second, select an inference framework matched to data characteristics: use Frequentist methods when identifiability is strong and data are rich; use Bayesian methods when identifiability is weak or data are limited. Third, interpret the empirical results in light of SI constraints, recognizing that no statistical method can uniquely recover structurally unidentifiable parameters.
For practitioners selecting between inference frameworks, we recommend: (1) Use Frequentist methods when minimizing point prediction error and computational cost is paramount, data quality is high, and models are structurally identifiable. (2) Use Bayesian methods when uncertainty quantification is essential, prior information is available, models contain unobserved states, or identifiability is limited. (3) Improve both methods by improving observational coverage, obtaining accurate initial condition estimates, and using SI-guided experimental design.
This comparative study advances our understanding of how statistical inference frameworks perform in biological modeling contexts and provides practical guidance for method selection. By aligning the inference approach with problem characteristics and grounding interpretation in identifiability theory, modelers can achieve reliable, reproducible, and scientifically defensible results that inform ecological management and public health decision-making.
Appendix A Supplementary Material
A.1 LV-Model (BFF) Histograms

### v2

@@ 6.3 GLM model for mpox epidemic in the USA
For the 2022 U.S. mpox epidemic, both methods reproduce the overall rise–fall pattern, indicating that the GLM captures the main outbreak trajectory under either inference framework (Figure 14). Parameter estimates are broadly consistent across methods, with modest differences in uncertainty widths (Table 20). For example, $p$ is tightly concentrated around $\sim 0.8$–0.85 for both methods, while $r$ differs more in its central estimate (BFF 1.95 vs. QDF 2.25), reflecting method-specific tradeoffs in fitting the growth and saturation phases.
Figure 14: Fitting visualization for the GLM model, with U.S. mpox data, using QDF and BFF approaches. Left QDF, and Right BFF.
Calibration-period metrics favor QDF on point-error measures (MAE 9.45 vs. 11.739; MSE 150.47 vs. 399.43; WIS 5.8 vs. 7.175), while both methods attain high PI coverage (100% for QDF vs. 97.8% for BFF; Table 21). The Bayesian fit converged cleanly ($\hat{R}=1$ for all GLM parameters; Table 22), supporting stability of the reported BFF intervals.
Table 20: Parameter estimates for the GLM model, with U.S. mpox data, using QDF and BFF approaches.
@@
The cross-method parameter comparison (Figure 15) and metric summary (Figure 16) reinforce these trends: parameter uncertainty is generally tight under both methods, and QDF attains lower MAE/MSE/WIS while maintaining near-nominal PI coverage.
Table 21: Performance metrics for the GLM model, with U.S. mpox data, using QDF and BFF approaches.
@@
Table 22: Convergence diagnostics for the GLM model, with U.S. mpox data, using BFF approache.
@@
$\hat{R}$ 1 1 1
Figure 15: Parameter comparison for the GLM model, with U.S. mpox data, using QDF and BFF approaches. We use a log scale for the x-axis.
Figure 16: Performance metrics comparison for the GLM model, with U.S. mpox data, using QDF and BFF approaches.
The parameter bounds used for applying the BFF and QDF for this model is provided in Table 23 and 24.
@@ 6.4 SEIUR model of COVID-19 cases in Spain during the first wave
For Spain’s first COVID-19 wave, both methods track the rise and decline in reported incidence (Figure 17), but the inferred uncertainties and error metrics differ in ways consistent with the SEIUR model’s limited identifiability under incidence-only observations (Section 5). Compared with QDF, BFF yields lower errors (MAE 294.57 vs. 352.30; MSE 210,754 vs. 311,054; WIS 211.30 vs. 250.46) with similar PI coverage (94% vs. 92%; Table 26), suggesting an advantage for Bayesian regularization in this partially observed, higher-dimensional setting.
Figure 17: Fitting visualization for the SEIUR model, with Spain’s first COVID-19 wave data, using QDF and BFF approaches. Left QDF, and Right BFF.
The parameter summaries (Table 25) differ most for $\rho$ and $q_{1}$, which are strongly coupled to unobserved compartments and therefore weakly constrained by incidence alone. This helps explain why the two frameworks can yield different central estimates despite producing similar fitted trajectories.
Convergence diagnostics indicate some remaining sampling difficulty for the Bayesian SEIUR fit (Table 27; several parameters have $\hat{R}>1.1$), which is expected in a coupled, weakly identified setting. Accordingly, we interpret SEIUR posterior summaries with appropriate caution.
@@
Table 25: Parameter estimates for the SEIUR model, with Spain’s first COVID-19 wave data, using QDF and BFF approaches.
Table 26: Performance metrics for the SEIUR model, with Spain’s first COVID-19 wave data, using QDF and BFF approaches.
@@
Table 27: Convergence diagnostics for the SEIUR model, with Spain’s first COVID-19 wave data, using BFF approach.
@@
$\hat{R}$ 1.18 1.22 1.17 1 1.07 1.21
The parameter comparison (Figure 18) highlights broader uncertainty for several rates under BFF, consistent with posterior uncertainty propagation under identifiability limitations.
Figure 18: Parameter comparison for the SEIUR model, with Spain’s first COVID-19 wave data, using BFF approach.
The metric summary (Figure 19) reinforces the table-based ranking: BFF achieves lower MAE/MSE/WIS while PI coverage remains close to nominal for both methods.
Figure 19: Performance metrics comparison for the SEIUR model, with Spain’s first COVID-19 wave data, using BFF approach.
The parameter bounds used for applying the BFF and QDF for this model is provided in Table 28 and 29.
Table 28: Parameter ranges for the SEIUR model, with Spain’s first COVID-19 wave data, using BFF approach.
@@
Table 29: Parameter ranges for the SEIUR model, with Spain’s first COVID-19 wave data, using QDF approach.
@@ 7 Discussion
This study compares Bayesian and Frequentist inference across ecological, clinical, and epidemiological models, with a focus on how data availability, model structure, and identifiability shape inference performance. Across all examples, we find that neither framework universally outperforms the other. Instead, performance depends critically on the degree of structural identifiability, the presence of latent states, and whether the primary goal is accurate point prediction or reliable uncertainty quantification. Bayesian methods, implemented via BayesianFitForecast (BFF), consistently provide stronger uncertainty estimates and diagnostic tools, while Frequentist methods, implemented via QuantDiffForecast (QDF), often achieve lower point prediction errors with substantially lower computational cost when identifiability is strong.
@@
For the Lotka–Volterra model, inference performance varied markedly across observation scenarios. When both prey and predator populations were observed, both methods produced stable and plausible parameter estimates. QDF yielded slightly lower point errors for predator dynamics, while BFF achieved better uncertainty calibration for prey. When only a single population was observed, QDF continued to provide sharper point predictions, whereas BFF maintained superior prediction interval coverage. Bayesian convergence diagnostics indicated stable posterior sampling even under partial observation, confirming the reliability of Bayesian inference in these settings.
Differences in parameter estimates across observation schemes reflect changes in practical identifiability. With full observation of both species, interaction and turnover parameters are more strongly informed by the data, reducing confounding but also exposing differences between inference philosophies. In this regime, Bayesian and optimization-based approaches may favor different trade-offs when fitting coupled dynamics, leading to larger discrepancies in some parameters. Under partial observation, the Lotka–Volterra system becomes more weakly identifiable, as multiple parameter combinations can reproduce similar marginal dynamics for the observed series while uncertainty is absorbed by the unobserved state. In this case, both methods are effectively constrained toward similar parameter regimes through regularization, bounds, or prior structure, which can make several parameter estimates appear closer despite increased uncertainty.
These empirical findings closely follow the predictions of the structural identifiability analysis. When both prey and predator are observed, all Lotka–Volterra parameters are structurally identifiable, whereas observing only one species leads to non-identifiability unless initial conditions are known. This limitation affects both Bayesian and Frequentist approaches equally, underscoring that observability places fundamental limits on inference quality regardless of statistical framework.
For the generalized logistic model applied to lung injury and mpox outbreaks, QDF consistently achieved better point forecast accuracy while maintaining similar or better uncertainty calibration. In the mpox case, QDF produced lower MAE, MSE, and WIS, while achieving near-perfect or perfect prediction interval coverage. Bayesian inference showed excellent convergence for all parameters, confirming well-sampled posteriors. Similar patterns were observed for lung injury. These results are explained by the strong structural identifiability of the generalized logistic model: the single-equation structure and direct observation of incident cases ensure that all parameters are identifiable, even when initial conditions are unknown. In this setting, optimization-based inference provides accurate estimates efficiently, while Bayesian inference adds computational overhead without clear gains in predictive accuracy.
A contrasting pattern emerges for the SEIUR model applied to COVID-19 data from Spain. Here, BFF outperformed QDF across all metrics, including point accuracy and uncertainty calibration. The SEIUR model includes latent compartments and exhibits limited structural identifiability when only reported cases are observed and initial conditions are unknown. Structural identifiability analysis shows that most parameters and state variables are unidentifiable in this setting, with only a small subset remaining identifiable. Under these conditions, Bayesian inference benefits from prior information, which regularizes the ill-posed inverse problem and stabilizes both parameter estimates and forecasts. In contrast, deterministic optimization must search a high-dimensional and weakly constrained parameter space, leading to less reliable results.
Taken together, these results show that structural identifiability sets hard limits on what can be learned from data. No statistical method can recover parameters that are structurally unidentifiable, and identifiability analysis should therefore precede model fitting. When identifiability is strong and data are rich, Frequentist methods are well suited for minimizing point forecast error and computational cost. When identifiability is weak, data are sparse or noisy, or models include latent states, Bayesian methods provide more reliable uncertainty quantification and stable inference.
Several methodological choices also influence these results. The use of different software environments reflects methodological requirements rather than implementation convenience: BFF relies on Stan-based Bayesian inference supported in R, while QDF depends on deterministic optimization and repeated numerical ODE solutions, for which MATLAB provides highly optimized solvers. Both approaches use identical model structures, data inputs, and evaluation metrics, ensuring that performance differences arise from inference methodology rather than computational platform.
In the practical identifiability analysis, initial conditions were treated as known and fixed to align with the structural identifiability setting and to isolate differences arising from inference methodology. Allowing initial conditions to be estimated would introduce additional confounding, particularly in partially observed systems, and would be expected to increase uncertainty and variability in parameter estimates. Both QDF and BFF support joint estimation of initial conditions, and examining the impact of this added uncertainty remains an important direction for future work.
This study intentionally focuses on low-dimensional models to allow transparent interpretation and the application of differential-algebraic structural identifiability methods. While this enables a clean methodological comparison, it also imposes limitations. We assumed normal error models, analyzed systems independently, and considered idealized identifiability conditions. For larger and more realistic models, practical identifiability tools based on Fisher information or sensitivity analysis may be necessary, especially under noisy or incomplete observations. Similarly, while Hamiltonian Monte Carlo performs well for moderate parameter dimensions, higher-dimensional models may require alternative sampling strategies such as adaptive Metropolis or DRAM.
@@
Future work could explore hybrid inference approaches that combine the computational efficiency of optimization with Bayesian uncertainty quantification, as well as extensions to higher-dimensional, spatial, stochastic, or agent-based models. Structural identifiability analysis could also be integrated more directly into experimental design, guiding data collection before inference is attempted. Incorporating machine learning approaches such as neural ordinary differential equations or physics-informed neural networks may provide complementary advantages when combined with identifiability-aware modeling.
Supplementary Material
7.1 LV-Model (BFF) Histograms